A pair of twins are 18.0 years old. When they leav
Answer #1

Problem 5)

The rest mass of the particle is,

m_0=135MeV/c^2

Here ,

\beta=\frac{v}{c}=\frac{12}{13}

So,

\gamma=\frac{1}{\sqrt{1-\beta^2}}=\frac{1}{\sqrt{1-12^2/13^2}}=2.6

a) So the energy of the particle is,

E=m\gamma c^2=135\times2.6MeV=351MeV

b) The momentum is p

E^2=p^2c^2+(m_0c^2)^2

p=\sqrt{E^2-(m_0c^2)^2}/c=\sqrt{351^2-135^2}MeV/c=324MeV/c

c) Kinetic energy is,

KE=E-m_0c^2

=(351-135) MeV=216MeV